OpenStudy (anonymous):

what is the answer to x2-5x+6=0

OpenStudy (callisto):

x^2-5x+6 = 0 x^2 - 2x - 3x + 6 = 0 x(x-2) - 3 (x-2) = 0 (x-3)(x-2) = 0 x-3 = 0 or x-2 = 0 Can you solve them?